Janet is 6 years older than 3 times the age of her son Michael.The difference between Michael's age and Janet's age is 28 years. How old is Janet and Michael?

Answer by tmwilliamlin168(27)

j=Janet
m=Michael
j=6+3m
j-m=28
6+2m=28
2m=22
m=11
j=39
Janet:39 years old
Michael: 11 years old